OpenStudy (anonymous):

What is the product of 5the square root of 3. • 4the square root of 15.? Simplify if possible.

OpenStudy (anonymous):

\[5\sqrt{3}\times4\sqrt{15}\]

OpenStudy (anonymous):

I think that's as simple as you can go

OpenStudy (anonymous):

\[\large\sqrt[5]{3}\times \sqrt[4]{15}?\]

OpenStudy (anonymous):

either way, that's is as simplified as it will get :)

OpenStudy (anonymous):

that*

OpenStudy (anonymous):

Ball is on your court. Now let's dance.!

OpenStudy (anonymous):

no its not simplified

OpenStudy (anonymous):

mofo it is simplified

OpenStudy (anonymous):

Well what is the product then?

OpenStudy (anonymous):

(5*4)(rad(3*15))

OpenStudy (anonymous):

I am to help. I said it is simplifies

OpenStudy (anonymous):

you can do more

OpenStudy (anonymous):

The only thing you can do is multiply the 4 and 5 to get 20

OpenStudy (anonymous):

\[(5\times4)\sqrt{3\times15}\]

OpenStudy (anonymous):

so would it be 20 sqrtrt 15?

OpenStudy (anonymous):

and you can extract a square root of 3 but that's it

OpenStudy (anonymous):

\(\sqrt{3\times 15}=\sqrt{3^2\times 5}=\sqrt{3^2}\times \sqrt{5}=3\sqrt{5}\)

OpenStudy (anonymous):

\[60 \sqrt{5}\]

OpenStudy (anonymous):

following nick you would get\[20\sqrt{45}\]which can be simplified to \[20\times3\sqrt{5}\]\[60\sqrt{5}\]

OpenStudy (anonymous):

Final answer boyah!

OpenStudy (anonymous):

Brohoof ?

OpenStudy (anonymous):

Could I also have help on finding the product of 6 square root of the quantity 12 times k. • 4kthe square root of 3.? Simplify if possible.

OpenStudy (anonymous):

do it the same way pound

OpenStudy (anonymous):

Im confused now though, would it be 144k sqrrt 2k?

OpenStudy (anonymous):

\[6\sqrt{12k}\times4k \sqrt{3}\rightarrow(6\times4k)\sqrt{12k \times3}\]

OpenStudy (anonymous):

(24k) sqrrt36k?

OpenStudy (anonymous):

ya now reduce

OpenStudy (anonymous):

Im confused as to how to reduce this

OpenStudy (anonymous):

72kthe square root of k. 144kthe square root of the quantity 2 times k. 72kthe square root of the quantity 2 times k. 144kthe square root of k. Those are my possible answers

OpenStudy (anonymous):

?